. Calculate the power conversion efficiency and maximum power (Pmax ) drawn from

a Si solar cell illuminated by sun light of intensity 100 mW/cm2 when Voc=0.7 V, Isc=3.5 A and 

FF=0.7. What is the ideal value of Pmax for solar cell


1
Expert's answer
2021-07-05T08:41:18-0400

"\eta=\frac {FF\cdot P}{P_0}=\frac {FF\cdot P}{UI}=\frac{0.7\cdot 0.1}{0.7\cdot 3.5}=0.029,"

"P_{max}=UI=0.7\cdot 3.5=24.5~W,"

"P_{id}=\frac{P_{max}}{FF}=\frac{24.5}{0.7}=35~W."